# 1]DAMO Academy, Alibaba Group

### v1

@@ 3.4 Training and Inference
Mask Alignment (Stage-1) General Mask Captioning 1.17M RefCOCO [71, 42], VideoRefer-Caption [75], DAM [32], Osprey-Caption [73], MDVP-Data [34], HC-STVG [56]
@@
0.54M VLM-3R-Data [17]
General VQA 0.74M LLaVA-OV-SI [27], LLaVA-Video [84], ShareGPT-4o-video [8], VideoGPT-plus [40], FineVideo [18], CinePile [52], ActivityNet [5], YouCook2 [86], LLaVA-SFT [36]
@@
General Segmentation 0.32M ADE20K [85], COCOStuff [6], Mapillary [45], PACO-LVIS [50], PASCAL-Part [9]
@@
General VQA 0.80M LLaVA-OV-SI [27], LLaVA-Video [84], ShareGPT-4o-video [8], VideoGPT-plus [40], FineVideo [18], CinePile [52], ActivityNet [5], YouCook2 [86], LLaVA-SFT [36]
@@ 4 Experiments
    GPT-4o    [46] 28.3 41.1 33.9 13.4 22.8 6.0 24.3 16.7 36.1 22.2
GPT-4.1 [46] 33.5 45.9 37.8 17.2 27.6 6.1 35.9 30.4 45.7 28.8
Seed1.5-VL [20] 34.7 52.1 42.8 8.2 27.7 4.3 32.9 19.1 27.9 26.1
Genimi-2.5 Pro [12] 45.5 64.0 52.7 9.3 36.7 8.1 47.0 29.9 69.3 37.8
@@
    VideoLLaMA3-7B    [76] 27.3 36.7 30.2 5.1 26.8 1.2 30.0 19.0 34.9 24.1
InternVL3-78B [87] 29.0 45.3 37.3 9.0 31.8 2.2 10.9 30.9 26.0 20.0
Qwen2.5-VL-72B [4] 36.4 54.2 44.7 11.3 24.8 7.2 27.2 22.9 83.7 27.4
@@ 4.2.1 Main Results
Tab. 2 presents the evaluation results of our RynnEC model and five categories of related MLLMs on the RynnEC-Bench. Although the RynnEC model contains only 7B parameters, it demonstrates robust embodied cognitive abilities, outperforming even the most advanced proprietary model, Gemini-2.5 Pro [12], by 10.7 points. Moreover, RynnEC achieves both balanced and superior performance across various tasks. For object cognition, RynnEC achieved a score of 61.4 and possesses the ability to both understand and segment objects. In terms of spatial cognition, RynnEC achieves a score of 54.5, which is 44.2% higher than that of Gemini-2.5 Pro. To support resource-constrained settings, we present a 2B-parameter RynnEC that delivers markedly lower inference latency while maintaining near-parity performance (<2<2 percentage points drop), enabling on-device deployment for embodied applications. In the following sections, we will introduce the performance of different types of MLLMs on RynnEC-Bench in detail.
@@
Qwen2.5-VL-72B [4] exhibits outstanding performance, achieving a score of 36.4 and surpassing GPT-4.1 [46]. This suggests that, in specialized capabilities such as embodied cognition, the gap between open-source and proprietary MLLMs has been significantly narrowed. Furthermore, we observe that Qwen2.5-VL and InternVL3 [87] demonstrate superior performance in positional relationship (PR) and distance perception tasks, respectively, even outperforming Gemini-2.5 Pro. Such pronounced differences in various aspects of spatial cognition may be attributed to the distribution of training data.
@@
These MLLMs are capable of accepting region masks as input, enabling more direct localization of target objects and facilitating finer-grained object perception. VideoRefer-VL3-7B [75] is a model fine-tuned from the base model VideoLLaMA3-7B [76]. As shown in Tab. 2, VideoRefer-VL3-7B consistently outperforms VideoLLaMA3-7B in both object cognition and spatial cognition tasks. This demonstrates that, in embodied scenarios, integrating mask understanding within the model is superior to explicit visual prompting.
@@
Recently, several studies have applied MLLMs to object segmentation tasks while retaining the original multimodal understanding capabilities of MLLMs. However, the best-performing model, RGA3-7B [59], achieves only 15.2 points on the object properties task. Although these MLLMs can still address some general video understanding tasks, their task generalization ability is significantly diminished following segmentation training. In contrast, our RynnEC model, which is specifically designed for embodied scenarios, maintains strong object and spatial understanding capabilities even after segmentation training.
@@
With the growing demand for highly generalizable cognitive abilities in the field of embodied intelligence, a number of studies have begun to develop MLLMs specifically tailored for embodied scenarios. A representative model is RoboBrain-2.0 [57], which achieves 24.2 even worse than general-purpose video models such as VideoLLaMA3-7B. There are two primary reasons for this: (1) Loss of object cognition: Embodied MLLMs typically emphasize spatial perception and task planning abilities, but tend to overlook the importance of detailed object understanding. (2) Lack of fine-grained perceptual capability: In egocentric videos, RoboBrain-2.0 demonstrates limited ability to interpret region-level features.
@@ 4.2.3 Spatial Cognition
  Qwen2.5-VL-7B [4] 35.9
InternVL3-8B [87] 42.1
GPT-4o [46] 43.6
Magma-8B [68] 12.7
Cosmos-Reason1-7B [3] 25.6
VeBrain-8B [37] 26.3
RoboBrain-7B-1.0 [25] 31.1
RoboBrain-7B-2.0 [57] 36.1
M2-Reasoning-7B [1] 42.3
@@
Figure 6: Performance on VSI-Bench [69]. Left: per-subtask comparison with VideoLLaMA3, the base model of our RynnEC. Right: overall comparison with generalist MLLMs and embodied MLLMs without explicit 3D encoding.
@@ 4.3 Generalization and Scalability
To investigate the generalizability of RynnEC, we conduct experiments on VSI-Bench [69], a purely textual spatial intelligence benchmark. As shown in Fig. 6, RynnEC-7B consistently surpasses VideoLLaMA3-7B across almost all capability dimensions. Notably, RynnEC is trained with a mask-centric spatial awareness paradigm, whereas all tasks in VSI-Bench involve purely textual spatial reasoning. This demonstrates that spatial awareness need not be constrained by the modality of representation, and spatial reasoning abilities can be effectively transferred across modalities. Further observation reveals substantial performance gains of RynnEC on the Route Planning task, despite this task not being included during training. This indicates that the navigation performance of embodied agents is currently constrained by foundational spatial perception capabilities, such as the understanding of direction, distance, and spatial relationships. Only with robust foundational spatial cognition can large embodied models achieve superior performance in high-level planning and decision-making tasks. Compared to other embodied MLLMs of comparable size, RynnEC-7B also achieves a leading score of 45.8.
@@
Figure 7: The example of RynnEC assisting robots in performing long-range tasks. The robot accomplishes the two designated tasks within the RoboTHOR simulator [14]. RynnEC facilitates the robot in achieving fine-grained environmental cognition throughout the task execution.
@@ 4.4 Embodied Application
Recently, some works [11, 63] have leveraged MLLMs as the "brain" to assist robots in planning tasks, perceiving environments, and making decisions. However, current MLLMs lack key capabilities such as spatial awareness, fine-grained perception, and instance localization, which restricts these applications to limited and simple tasks. As illustrated in Fig. 7, RynnEC demonstrates the potential to assist robots in accomplishing long-horizon tasks within complex environments. From two real-time tasks performed by the robot equipped with RynnEC, we observe the following roles that RynnEC plays in task execution: (1) Fine-grained object localization and understanding enable robots to more quickly identify target objects and assess their states; (2) Direction and distance perception of targets improves navigation efficiency and precision; (3) Spatial scale estimation empowers robots to perform more delicate manipulations; (4) Counting ability facilitates the completion of tasks requiring mathematical reasoning. It is important to emphasize that the role of RynnEC in embodied tasks is far from limited to these examples. We hope that more researchers will integrate RynnEC models into robotic systems across a wide range of tasks, thereby advancing embodied intelligence toward more valuable real-world applications.
@@ 5 Conclusion and Future Works
  • Unified Perception and Planning Framework: Recent studies [57] have started to explore training unified embodied intelligence models that combine perception and planning. However, these approaches are limited in their ability to facilitate fine-grained, video-based visual interactions. In the future, we aim to endow RynnEC with more flexible planning abilities and integrate it with VLA models to form a closed-loop embodied system.

### v2

@@ 3.4 Training and Inference
Mask Alignment (Stage-1) General Mask Captioning 1.17M RefCOCO Yu et al. [2016], Mao et al. [2016], VideoRefer-Caption Yuan et al. [2025c], DAM Lian et al. [2025], Osprey-Caption Yuan et al. [2024], MDVP-Data Lin et al. [2024], HC-STVG Tang et al. [2021]
@@
0.54M VLM-3R-Data Fan et al. [2025]
General VQA 0.74M LLaVA-OV-SI Li et al. [2024a], LLaVA-Video Zhang et al. [2024e], ShareGPT-4o-video Chen et al. [2024b], VideoGPT-plus Maaz et al. [2024], FineVideo Farré et al. [2024], CinePile Rawal et al. [2024], ActivityNet Caba Heilbron et al. [2015], YouCook2 Zhou et al. [2018], LLaVA-SFT Liu et al. [2023]
@@
General Segmentation 0.32M ADE20K Zhou et al. [2017], COCOStuff Caesar et al. [2018], Mapillary Neuhold et al. [2017], PACO-LVIS Ramanathan et al. [2023], PASCAL-Part Chen et al. [2014]
@@
General VQA 0.80M LLaVA-OV-SI Li et al. [2024a], LLaVA-Video Zhang et al. [2024e], ShareGPT-4o-video Chen et al. [2024b], VideoGPT-plus Maaz et al. [2024], FineVideo Farré et al. [2024], CinePile Rawal et al. [2024], ActivityNet Caba Heilbron et al. [2015], YouCook2 Zhou et al. [2018], LLaVA-SFT Liu et al. [2023]
@@ 4 Experiments
    GPT-4o        OpenAI et al. [2024] 28.3 41.1 33.9 13.4 22.8 6.0 24.3 16.7 36.1 22.2
GPT-4.1 OpenAI et al. [2024] 33.5 45.9 37.8 17.2 27.6 6.1 35.9 30.4 45.7 28.8
Seed1.5-VL Guo et al. [2025] 34.7 52.1 42.8 8.2 27.7 4.3 32.9 19.1 27.9 26.1
Genimi-2.5 Pro Comanici et al. [2025] 45.5 64.0 52.7 9.3 36.7 8.1 47.0 29.9 69.3 37.8
@@
    VideoLLaMA3-7B        Zhang et al. [2025a] 27.3 36.7 30.2 5.1 26.8 1.2 30.0 19.0 34.9 24.1
InternVL3-78B Zhu et al. [2025] 29.0 45.3 37.3 9.0 31.8 2.2 10.9 30.9 26.0 20.0
Qwen2.5-VL-72B Bai et al. [2025] 36.4 54.2 44.7 11.3 24.8 7.2 27.2 22.9 83.7 27.4
@@ 4.2.1 Main Results
Tab. 2 presents the evaluation results of our RynnEC model and five categories of related MLLMs on the RynnEC-Bench. Although the RynnEC model contains only 7B parameters, it demonstrates robust embodied cognitive abilities, outperforming even the most advanced proprietary model, Gemini-2.5 Pro Comanici et al. [2025], by 10.7 points. Moreover, RynnEC achieves both balanced and superior performance across various tasks. For object cognition, RynnEC achieved a score of 61.4 and possesses the ability to both understand and segment objects. In terms of spatial cognition, RynnEC achieves a score of 54.5, which is 44.2% higher than that of Gemini-2.5 Pro. To support resource-constrained settings, we present a 2B-parameter RynnEC that delivers markedly lower inference latency while maintaining near-parity performance (<2<2 percentage points drop), enabling on-device deployment for embodied applications. In the following sections, we will introduce the performance of different types of MLLMs on RynnEC-Bench in detail.
@@
Qwen2.5-VL-72B Bai et al. [2025] exhibits outstanding performance, achieving a score of 36.4 and surpassing GPT-4.1 OpenAI et al. [2024]. This suggests that, in specialized capabilities such as embodied cognition, the gap between open-source and proprietary MLLMs has been significantly narrowed. Furthermore, we observe that Qwen2.5-VL and InternVL3 Zhu et al. [2025] demonstrate superior performance in positional relationship (PR) and distance perception tasks, respectively, even outperforming Gemini-2.5 Pro. Such pronounced differences in various aspects of spatial cognition may be attributed to the distribution of training data.
@@
These MLLMs are capable of accepting region masks as input, enabling more direct localization of target objects and facilitating finer-grained object perception. VideoRefer-VL3-7B Yuan et al. [2025c] is a model fine-tuned from the base model VideoLLaMA3-7B Zhang et al. [2025a]. As shown in Tab. 2, VideoRefer-VL3-7B consistently outperforms VideoLLaMA3-7B in both object cognition and spatial cognition tasks. This demonstrates that, in embodied scenarios, integrating mask understanding within the model is superior to explicit visual prompting.
@@
Recently, several studies have applied MLLMs to object segmentation tasks while retaining the original multimodal understanding capabilities of MLLMs. However, the best-performing model, RGA3-7B Wang et al. [2025], achieves only 15.2 points on the object properties task. Although these MLLMs can still address some general video understanding tasks, their task generalization ability is significantly diminished following segmentation training. In contrast, our RynnEC model, which is specifically designed for embodied scenarios, maintains strong object and spatial understanding capabilities even after segmentation training.
@@
With the growing demand for highly generalizable cognitive abilities in the field of embodied intelligence, a number of studies have begun to develop MLLMs specifically tailored for embodied scenarios. A representative model is RoboBrain-2.0 Team et al. [2025], which achieves 24.2 even worse than general-purpose video models such as VideoLLaMA3-7B. There are two primary reasons for this: (1) Loss of object cognition: Embodied MLLMs typically emphasize spatial perception and task planning abilities, but tend to overlook the importance of detailed object understanding. (2) Lack of fine-grained perceptual capability: In egocentric videos, RoboBrain-2.0 demonstrates limited ability to interpret region-level features.
@@ 4.2.3 Spatial Cognition
  Qwen2.5-VL-7B Bai et al. [2025] 35.9
InternVL3-8B Zhu et al. [2025] 42.1
GPT-4o OpenAI et al. [2024] 43.6
Magma-8B Yang et al. [2025b] 12.7
Cosmos-Reason1-7B Azzolini et al. [2025] 25.6
VeBrain-8B Luo et al. [2025] 26.3
RoboBrain-7B-1.0 Ji et al. [2025] 31.1
RoboBrain-7B-2.0 Team et al. [2025] 36.1
M2-Reasoning-7B AI et al. [2025] 42.3
ViLaSR Wu et al. [2025b] 45.4
@@
Figure 6: Performance on VSI-Bench Yang et al. [2025c]. Left: per-subtask comparison with VideoLLaMA3, the base model of our RynnEC. Right: overall comparison with generalist MLLMs and embodied MLLMs without explicit 3D encoding.
@@ 4.3 Generalization and Scalability
To investigate the generalizability of RynnEC, we conduct experiments on VSI-Bench Yang et al. [2025c], a purely textual spatial intelligence benchmark. As shown in Fig. 6, RynnEC-7B consistently surpasses VideoLLaMA3-7B across almost all capability dimensions. Notably, RynnEC is trained with a mask-centric spatial awareness paradigm, whereas all tasks in VSI-Bench involve purely textual spatial reasoning. This demonstrates that spatial awareness need not be constrained by the modality of representation, and spatial reasoning abilities can be effectively transferred across modalities. Further observation reveals substantial performance gains of RynnEC on the Route Planning task, despite this task not being included during training. This indicates that the navigation performance of embodied agents is currently constrained by foundational spatial perception capabilities, such as the understanding of direction, distance, and spatial relationships. Only with robust foundational spatial cognition can large embodied models achieve superior performance in high-level planning and decision-making tasks. Compared to other embodied MLLMs of comparable size, RynnEC-7B also achieves a leading score of 45.8.
@@
Figure 7: The example of RynnEC assisting robots in performing long-range tasks. The robot accomplishes the two designated tasks within the RoboTHOR simulator Deitke et al. [2020]. RynnEC facilitates the robot in achieving fine-grained environmental cognition throughout the task execution.
@@ 4.4 Embodied Application
Recently, some works Cheng et al. [2024b], Xiong et al. [2024] have leveraged MLLMs as the "brain" to assist robots in planning tasks, perceiving environments, and making decisions. However, current MLLMs lack key capabilities such as spatial awareness, fine-grained perception, and instance localization, which restricts these applications to limited and simple tasks. As illustrated in Fig. 7, RynnEC demonstrates the potential to assist robots in accomplishing long-horizon tasks within complex environments. From two real-time tasks performed by the robot equipped with RynnEC, we observe the following roles that RynnEC plays in task execution: (1) Fine-grained object localization and understanding enable robots to more quickly identify target objects and assess their states; (2) Direction and distance perception of targets improves navigation efficiency and precision; (3) Spatial scale estimation empowers robots to perform more delicate manipulations; (4) Counting ability facilitates the completion of tasks requiring mathematical reasoning. It is important to emphasize that the role of RynnEC in embodied tasks is far from limited to these examples. We hope that more researchers will integrate RynnEC models into robotic systems across a wide range of tasks, thereby advancing embodied intelligence toward more valuable real-world applications.
@@ 5 Conclusion and Future Works
  • Unified Perception and Planning Framework: Recent studies Team et al. [2025] have started to explore training unified embodied intelligence models that combine perception and planning. However, these approaches are limited in their ability to facilitate fine-grained, video-based visual interactions. In the future, we aim to endow RynnEC with more flexible planning abilities and integrate it with VLA models to form a closed-loop embodied system.